OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5

OpenStudy (anonymous):

@falseyy can you help?

OpenStudy (anonymous):

-4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

jigglypuff314 (jigglypuff314):

I would suggest first solving the system of equations 4x - y = -11 2x + 3y = 5 for x and y...

OpenStudy (anonymous):

yea thats what i'm doing now

jigglypuff314 (jigglypuff314):

okay, what have you got for x so far?

OpenStudy (anonymous):

well for x (i think it's wrong) i got 6.3 repeating.

OpenStudy (anonymous):

can you please just help me quickly I really need this answer!!!

OpenStudy (anonymous):

@sweetburger can you please help?

jigglypuff314 (jigglypuff314):

hmm... here is what I did (I used substitution method) 4x - y = -11 --> y = 4x + 11 2x + 3y = 5 = 2x + 3(4x + 11) = 5 distribute 2x + 12x + 33 = 5 add like terms and subtract both sides by 33 14x = -28 <-- (shortcut: which option has that?)

OpenStudy (anonymous):

thank you!!!!!!!! Can you help me on one other?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

jigglypuff314 (jigglypuff314):

2x + y = 1 --> y = -2x + 1 4x + 2y = -1 = 4x + 2(-2x + 1) = -1 distribute 4x - 4x + 2 = -1 add like terms 2 = -1 2 does not equal -1 therefore no solution but be careful because if it was something like 2=2 then it would be infinite number of solutions

OpenStudy (anonymous):

okay thanks soo much

jigglypuff314 (jigglypuff314):

glad I could help :)